OpenStudy (anonymous):

Write an inequality that represents the mileage of any route between Portland and Tucson?

OpenStudy (hari5719):

dont know sry

OpenStudy (anonymous):

ohh lol ok

OpenStudy (phebe):

@Wolfboy

OpenStudy (anonymous):

@satellite73 are u going to help

OpenStudy (anonymous):

they are all greater than or equal to the miles of the direct flight

OpenStudy (anonymous):

if i am reading it correctly the direct flight is 973 miles so you could say for example \[x\geq 973\]

OpenStudy (anonymous):

ok thanx soo much can u help wit one more

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thanx soo much

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

Question 2 Your travel agent is making plans for you to go from San Diego to Seattle. A direct flight is not available. Option A consists of flights from San Diego to Boise to Seattle. Option B consists of flights from San Diego to Las Vegas to Seattle. Write an inequality to compare the distances of these two options.

OpenStudy (anonymous):

one is longer than the other i bet you can figure out which is which

OpenStudy (anonymous):

yes i can San Diego , Boise to sealttle right?

OpenStudy (anonymous):

i dunno i lost the page what number is it? page number i mean

OpenStudy (anonymous):

169

OpenStudy (anonymous):

lol u must have been looking at the other pages right

OpenStudy (anonymous):

225+735 is one right? yes i was

OpenStudy (anonymous):

u sure i can use that

OpenStudy (anonymous):

and 651+347

OpenStudy (anonymous):

no hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait nvm start all over please ur right

OpenStudy (anonymous):

Option A consists of flights from San Diego to Boise to Seattle. \[651+347=998\]

OpenStudy (anonymous):

yes it odes ohhh like that , ok

OpenStudy (anonymous):

Option B consists of flights from San Diego to Las Vegas to Seattle\[225+753=978\]

OpenStudy (anonymous):

i will let you figure out which one is larger

OpenStudy (anonymous):

Obtion is longer